OpenStudy (anonymous):

Find the equation of a line given the point and slope below. Arrange your answer in the form y = mx + b, where b is the constant. (5, 5) m = -3

OpenStudy (jdoe0001):

\(\bf \begin{array}{lllll} &x_1&y_1\\ &(5\quad ,&5) \end{array} \\\quad \\ slope = m= -3 \\ \quad \\ y-y_1=m(x-x_1)\quad \textit{plug in your values and solve for "y"} \)